OpenStudy (anonymous):

Integrate 9^x using the Right Riemann Sum with 4 intervals from x=0 to x=2

OpenStudy (anonymous):

I got to \[\frac{ 1 }{ 2 }\sum_{n = 1}^{4} 3^{n}\] What do I do now?

geerky42 (geerky42):

How did you end up with \(3^n\)?

geerky42 (geerky42):

I think it should be \[\frac{ 1 }{ 2 }\sum_{n = 1}^{4} 9^{n/2}\]

geerky42 (geerky42):

Since you just go from n=0 to 4, you can just calculate directly.

geerky42 (geerky42):

\[\frac{ 1 }{ 2 }\sum_{n = 1}^{4} 9^{n/2} = \dfrac{1}{2}(9^{0.5}+9^1+9^{1.5}+9^2)\]

geerky42 (geerky42):

ohh now I see how you have \(3^n\)

geerky42 (geerky42):

either way, just calculate directly.

OpenStudy (anonymous):

thanks!!

geerky42 (geerky42):

silly my brain... Welcome!